40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 12 01:09:56 2016, gautam, Update, LSC, Recovering DRMI locking 
    Reply  Wed Nov 16 03:10:01 2016, gautam, Update, LSC, DRMI locked on 1f and 3f signals 
       Reply  Wed Nov 16 08:14:43 2016, Steve, Update, LSC, DRMI locked on 1f and 3f signals 5hrs.png
       Reply  Mon Nov 21 14:02:32 2016, gautam, Update, LSC, DRMI locked on 3f signals, arms held on ALS DRMIArms_Nov20.pdf
    Reply  Wed Nov 23 16:21:02 2016, gautam, Update, LSC, ITMY UL glitches are back ITMYproblemsBack.png
       Reply  Mon Nov 28 10:27:13 2016, gautam, Update, SUS, ITMY UL glitches are back ITMY_satboxSpectra.pdfITMY_UL_testerBox.png
       Reply  Tue Nov 29 11:07:37 2016, Steve, Update, LSC, ITMY UL glitches are back gliching400d.png
          Reply  Wed Nov 30 01:47:56 2016, gautam, Update, LSC, Suspension woes IMCwoes.png
             Reply  Wed Nov 30 17:08:56 2016, gautam, Update, LSC, Binary output breakout box removed 
                Reply  Thu Dec 1 02:19:13 2016, gautam, Update, LSC, Binary output breakout box restored MC1_glitch_fast.pngITMY_testerGlitch.pngITMYsatbox.pngULcomparison.pdf
                   Reply  Thu Dec 1 08:02:57 2016, Steve, Update, LSC, glitching ITMY_UL_LL ITMY_UL_LL.png
                      Reply  Fri Dec 2 11:56:42 2016, gautam, Update, LSC, MC1 LEMO jiggled IMG_3474.JPGMC1_LR_dead.png
                         Reply  Mon Dec 5 15:05:37 2016, gautam, Update, LSC, MC1 glitches are back Mc1glitches.png
                      Reply  Thu Dec 8 10:13:43 2016, Steve, Update, LSC, glitching ITMY_UL has a history glitching__ITMY-UL_2007.png
Message ID: 12643     Entry time: Mon Nov 28 10:27:13 2016     In reply to: 12638
Author: gautam 
Type: Update 
Category: SUS 
Subject: ITMY UL glitches are back 

I left the tester box plugged in from Thursday night to Sunday afternoon, and in this period, the glitches still appeared in (and only in) the UL channel.

So yesterday evening, I pulled the Sat. Box. out and checked the DC voltages at various points in the circuit using a DMM, including the output of the high current buffer that supplies the drive current to the shadow sensor LEDs. When we had similar behaviour in the PRM box, this kind of analysis immediately identified the faulty component as the high current buffer IC (LM6321M) in the bad channel, but everything seems in order for the ITMY box. 

I then checked the Satellite Amplifier Termination Board, which basically just adds 100ohm series resistors to the output of the PD readout, and all the resistors seem fine, the piece of insulating material affixed to the bottom of this board is also intact. I then used the SR785 in AC coupled mode to look at the high frequency spectrum at the same points I checked the DC voltages with the DMM (namely the drive voltage to the LEDs, and the PD readout voltages on the PCB as well as on the pins of the connector on the outside of the box after the termination board (leading to the DAQ), and nothing sticks out here in the UL channel either. Of course it could be that the glitches are intermittent, and during my tests they just weren't there...

I am hesitant to start pulling out ICs and replacing them without any obvious signs of failure from them, but I am out of debugging ideas...


One possibility is that the problem lies upstream of the Sat. Box - perhaps the UL channel in the Suspension PD Whitening and Interface Board is faulty. To test, I have now hooked up ITMY Sat. Box. + tester box to the signal chain of ETMY. If I can get the other tester box back from Ben, I will plug in the ETMY sat. box. + tester to the ITMY signal chain. This should tell us something...

Attachment 1: ITMY_satboxSpectra.pdf  150 kB  | Hide | Hide all
ITMY_satboxSpectra.pdf
ELOG V3.1.3-